# USTC-ICTS/PCFT-24-24

### v1

@@ III.1 Dipolar (charge-swapping) boson star
We find that if the initial distance is too small, or the solution too compact, an apparent horizon is formed and the scalar field configuration results in gravitational collapse within $t\sim 10^{3}$. For a fixed value of $\omega$, we observe that as the distance $d$ increases, a greater amount of energy is radiated away in the initial relaxation period. Nevertheless, if the stars in Table 1 are placed sufficiently far away from each other initially, they can attract and relax to form a charge-swapping configuration. A typical scenario of a charge-swapping boson star forming is shown in Fig. 4. We see that in this case the initial configuration is actually relatively close to the charge-swapping bound state. We see that the U(1) charge in the $z>0$ half-space, $Q_{\text{up}}$, starts oscillating with an amplitude slightly smaller that the initial one, which is due to a small portion of the initial configuration being ejected shortly after the simulation begins. In the top panel of Fig. 4, we show half of a charge-swapping period, with the first plot and the last plot showing that the positive and negative charges have been completely inverted during this half period.
For $\omega=0.98$, we find that below the value $d=24$, which is the one showed in Fig. 4, the next case, which is $d=18$, also forms a regular horizonless final configuration, but the configuration with $d=12$ forms a black hole. So according to the mapping of the parameter space we use, the minimum distance to form a boson-anti boson regular system is $d_{\min}=18$. For $\omega=0.97$, which is both more massive and compact, we find a bigger minimum distance $d_{\min}=24$ is required. The trend continues with the $0.96$ and $0.95$ cases as displayed in the last column of Table 1.
@@
One can verify that the difference between the frequencies of the real and imaginary component of the scalar field is the origin of the dominant oscillation mode of the charge distribution. To see this, we make a spectral decomposition of the scalar field at a point outside the origin—at the origin, the imaginary part of $\phi$ is zero by construction. We see from Fig. 6 that the scalar field is practically monochromatic and the difference between the maximum of the real and imaginary parts gives $\omega_{R}-\omega_{I}=0.020$ (see the caption of Fig. 6 for more details), which is consistent with the dominant peak of $Q^{\rm up}$. In addition, we also find a second peak which is located around the first odd multiple of the base frequency, analogously to what was found in Xie et al. (2021) and consistent with the fact that a scalar potential has a $\mathbb{Z}_{2}$ symmetry. On the other hand, the small oscillations of $Q^{\rm up}$ that we already discussed corresponding the the peak in the bottom right panel of Fig. 5, coincide very well with $\omega_{R}+\omega_{I}=1.898$ and can be shown to be produced by the next order perturbation term on the charge density $\propto j^{0}$, after the term with frequency $\omega_{R}-\omega_{I}$.
@@ III.2 Higher multipolar boson stars
We find that none of the first five types of spherical stars in Table 1 leads to charge-swapping configurations for the same range of $d$ as in the dipole case or even if extending it to $d=48$. Black holes are often formed instead. In contrast, the last two types of stars with frequencies of $0.99$ and $0.995$ are capable of forming non-trivial configurations that are everywhere regular. Stationary boson stars with these frequencies are very dilute, as can be seen from their values of $\mathcal{C}$ and radius $R_{99}$ in Table 1. We have tried three $d$ distances for these two $\omega$’s respectively: 48, 42 and 36, and find that for both $\omega$’s no apparent horizon is formed for $d=42$ and 48, but for $d=36$ (and also smaller values), the initial clump collapses to a black hole. In Fig. 8, we can see the collapse of the lapse for one of the quadrupole cases together with the corresponding dipolar case which does not collapse.
@@
Configurations that avoid collapse exhibit quadrupolar charge-swapping structures. Despite differing shapes, these structures show quantitative similarities to the dipolar case, following similar correlations between initial parameters $\omega$ and $d$, and the final solution properties. However, their lifetimes differ markedly from self-interacting quadrupoles without gravity (i.e., quadrupolar Q-balls) Xie et al. (2021), with all cases persisting for at least $t=5000T_{0}$. In Fig. 9, snapshots of charge density illustrate typical behavior, comparing initial star distributions at $t=0$ with those within half a charge-swapping period later in the simulation. The bottom panel of Fig. 9 shows the total charge associated with one blob by integrating $n_{\mu}j^{\mu}$ over the $z>|x|$ region.
@@ IV.1 Sextic potential
until the turning point $\omega_{\rm min}$. This leads to one or two stable branches according to the value of $\alpha_{G}$; see Fig. 1. These observations are consistent with the conclusions drawn from a stability analysis of such solutions using catastrophe theory Tamaki and Sakai (2010); Kleihaus et al. (2012). After the turning point, all solutions are unstable regardless of the sign of $dM/d\omega$, some of them migrating towards a stable boson star and others dissipating to infinity or collapsing to a black hole depending on the binding energy, which is similar to the free field and the quartic self-interaction cases previously reported in the literature Seidel and Suen (1990); Balakrishna et al. (1998).
Compared to the free field case, the self-interacting solutions are qualitatively very different when the gravitational coupling is small, where they are close to their Q-ball counterparts. Nevertheless, these cases can give rise to very compact solutions (as $f(r=0)$ increases) that differ significantly from the Q-ball solutions and achieve very high compactness, making them different from both the mini-boson stars and the corresponding Q-balls; see Fig. 3. With this in mind, we have chosen the stars in Table 3 to prepare the charge-swapping configurations. In particular, we restrict our study to superpositions of single star solutions that are stable on their own. Of course, this does not guarantee that the complex boson star solution will be obtained, nor does it ensure that the final solution will not be a black hole. Since we have fixed $\mathfrak{g}=1/2$ as our fiducial model, we only have two degrees of freedom to construct the initial data: $\omega$ and $d$.
@@
We shall explore the dipole superposition with initial separations $d=\{8,10,12,14,16\}$. Before proceeding with the boson star analysis, as a sanity check, we first re-produce the 3+1D results of charge-swapping Q-balls in Ref. Xie et al. (2021). In Fig. 10, we see that the longest lifetime is given by $\omega=0.8$ and $d=14$, consistent with that of Ref. Xie et al. (2021), among other results. These flat-space solutions have smaller radii than the free-field configurations with gravity discussed in Section III. Nevertheless, for an easy comparison, we keep the same numerical setting as in the dipolar free field case with gravity.
@@
In essence, there are four stages of the charge-swapping configurations: An initial relaxation stage, which ends around $t=500T_{0}$ for the case presented in Fig. 10; a plateau stage of charge-swapping, ending around $t=1000T_{0}$; a fast decay stage in which $Q^{\rm up}$ drops abruptly to zero and the energy density also decreases but towards a finite spherically symmetric state; a long-lived stage of the (highly perturbed) oscillaton Seidel and Suen (1991) (the oscillon in the absence of gravity).
Next, we switch on $\alpha_{G}$ to investigate the effect of introducing the gravitational interaction. To monitor the evolution in this case, although not perfect as it is now not really conserved, we will still use the $E$ integral as a proxy. In all the cases presented in this section, we observe that the envelope of the Noether charge in the upper hemisphere always follows a similar qualitative behavior to that of the $E(t)$ curve. In Fig. 11, we show the cases with frequency $\omega=0.8$ and gravitational coupling $0.0018\pi\alpha_{G}=$ 0.001 and $0.01$. This corresponds to the two first rows in Table 3. For $0.0018\pi\alpha_{G}=$ 0.001, $\omega=0.8$ is the point where the spherical boson stars start to deviate from the Q-ball counterparts (see Fig. 1). The same is also noticeable in the top panel of Fig. 11 for the charge-swapping configurations, where we see that the lifetimes are slightly shortened in the presence of gravity for all the $d$’s listed. For a smaller $\alpha_{G}$, the system basically reduce to the Q-ball case. For $0.018\pi\alpha_{G}=$ 0.01 in the bottom panel, interestingly, we see that the differences between the $E$ curves are negligible for different $d$’s, due to the significantly reduced lifetimes. This means that the gravitational interaction is already dominating for $0.018\pi\alpha_{G}=$ 0.01, and the stable end state corresponds to an oscillaton as expected.
@@
The evolution of $E$ for $\omega=0.75$ and $0.018\pi\alpha_{G}=$ 0.01 for three different $d$’s together with the total Noether charge in the $z>0$ region for $d=16$ is shown in Fig. 12. We can see how the charge is radiated away in the final stage. More interesting is the fact that for these cases, as well as for the cases in Fig. 13, the configuration transits through two charge-swapping plateaus with different $E$ and $Q^{\rm up}$ before migrating to the oscillaton configuration. This resembles the cascading of the energy for excited oscillons Wang et al. (2023). However, at the level of the quantities we have observed, a directi connection between the two phenomena has yet to be established.
Figure 12: Varying $d$ for fixed $\omega=0.75$. In the right panel we show the Noether charge in the $z>0$ region, $Q^{\rm up}$, for the $d=16$ case.
@@ IV.2 Logarithmic potential
Also, similar to the charge-swapping Q-ball case with the logarithmic potential, there is no relaxation process that radiates away excess of the scalar field, as can be confirmed in Fig. 15. Also, the charge-swapping process begins very early in the simulations. The latter is reflected in the fact that the charge-swapping period $T_{\rm swap}$ is clearly defined from the beginning of the evolution, as can be seen in Fig. 16: $T_{\rm swap}$ is defined as twice the time elapsed between two contiguous zero points of the Noether charge in the upper half space $Q^{\rm up}$.
@@ V Anisotropic distribution of satellite galaxies
$(\Psi_{\rm w},U,x^{i},t)\rightarrow(\Lambda^{2}\Psi_{\rm w},\Lambda^{2}U,\Lambda^{-1}x^{i},\Lambda^{-2}t)\,.$ (68)
This allows to obtain the full sequence of Newtonian boson stars once a single solution has been found, unlike the relativistic analogue, where the family of solutions, say, the one presented in Fig. 1, must be constructed numerically solving the equations at each point. This rescaling is independent of the rescaling used to obtain dimensionless quantities, so to recover the physical system, as in the relativistic case, we must choose a value of $\mu$ as explained in Section II.3. We have previously obtained the metrics in the full general relativistic formulation. For this application, it is sufficient to extract the non-relativistic limit from Eq. (66) and Eq. (67).
We choose to perform the analysis at a certain representative space slice with $t=4000T_{0}$ of the dipolar configuration with $\omega=0.98$ well within the charge-swapping stage. However, the results are insensitive to the specific time period chosen within this stage. To extract the 2D data at such time we use the Kuibit tool Bozzola (2021). Based on the consistency statistical tests performed in Solís-López et al. (2021), we choose the values of $\mu$ and $\Lambda$, which completely fix the scale of the system, to be such that the size of the configuration, which has an equatorial radius of about 16 units, corresponds to a physical size of 300 kpc (radius encompassing the 11 classical satellites of the Milky Way Pawlowski and Kroupa (2020)), together with the condition that the circular velocity of the stars in the galactic disk at a point 30kpc away from the galactic center is 100 km/s, implying similar contributions from the enclosed mass at radius r=30kpc of dark matter and baryonic matter. This two conditions lead to the values $\mu\hbar\sim 10^{-25}\text{eV}$ and $\Lambda=0.01$. The precise determination of these two parameters given the rotation curve of the Milky Way and a model for the galactic bulge and disc deserves a statistical analysis which is beyond the scope of this paper.
@@
Figure 19: Angular distribution of the angular momenta of $10^{4}$ “particles” for the fiducial dipolar charge-swapping configuration. The initial distribution is isotropic, and it quickly relaxes to a configuration where the tangent of the orbital poles are oriented towards the galactic plane $\theta=\pi/2$ We only include particles within $\Lambda\mu R\in(0.16,16)$ which takes into account “particles” with $R>30$ kpc, the size of the Milky Way stellar disc.
@@
after restoring units. Meaning that the angular distribution of the angular momenta is already oriented toward the galactic pole after 1 Gyr (as can be seen in Fig. 19) and that the charge-swapping configurations live for much longer than is required for the galactic halos in the Universe.
@@
Figure 21: “Particles” in a disc of radius $30$ kpc and thickness $3$ kpc moving in a (Newtonian) gravitational potential of the fiducial dipolar configuration (see Fig. 18). The side of the cube is of length $60$ kpc.

### v2

@@ III.1 Dipolar (charge-swapping) boson star
We find that if the initial distance is too small, or the solution too compact, an apparent horizon is formed and the scalar field configuration results in gravitational collapse within $t\sim 10^{3}$. For a fixed value of $\omega$, we observe that as the distance $d$ increases, a greater amount of energy is radiated away in the initial relaxation period. Nevertheless, if the stars in Table 1 are placed sufficiently far away from each other initially, they can attract and relax to form a charge-swapping configuration.
A typical scenario of a charge-swapping boson star forming is shown in Fig. 4. The first observation is that, unlike their flat spacetime counterparts with a polynomial potential Xie et al. (2021), there is no indication of decay or a subsequent final oscillaton stage with total $Q=0$. For all configurations where a charge-swapping boson star successfully forms in the free field case, the evolution exhibits only two qualitatively distinct stages: an initial relaxation phase, followed by a charge-swapping plateau.
The first stage is depicted in the left inset of Fig. 4. In this specific case, we consider initial boson stars with a radius of $R_{99}=27.3$ , initially separated by a distance of $d=24$ Consistent with these parameters, the charge swap begins promptly after $t=0$, indicating that the initial configuration is relatively close to the charge-swapping bound state. We observe that the U(1) charge in the $z>0$ half-space, $Q^{\text{up}}$ , starts oscillating with an amplitude slightly smaller than the initial one. This reduction is due to a small portion of the scalar field being ejected shortly after the simulation begins. Although, in this example, the stars are already very close to each other at $t=0$, it is important to say, as we will discuss later, that in some cases, positioning the stars farther apart is crucial for achieving a charge-swapping pattern. In these cases we will refer to the early part of the initial relaxation stage, when the stars meet and start interacting both gravitationally and through scalar field interactions leading to the interchange of Noether charge, as the merger phase.
In the top panel of Fig. 4, we show half of a charge-swapping period, with the first plot and the last plot showing that the positive and negative charges have been completely inverted during this half period.
For $\omega=0.98$, we find that below the value $d=24$, which is the one showed in Fig. 4, the next case, which is $d=18$, also forms a regular horizonless final configuration, but the configuration with $d=12$ forms a black hole. So according to the mapping of the parameter space we use, the minimum distance to form a boson-anti boson regular system is $d_{\min}=18$. For $\omega=0.97$, which is both more massive and compact, we find a bigger minimum distance $d_{\min}=24$ is required. The trend continues with the $0.96$ and $0.95$ cases as displayed in the last column of Table 1. Of course, the configurations with smaller $\omega$ are initially smaller and, as they require a larger initial distance to avoid collapse, the merger of the stars is delayed. For these cases, and before the merger, the charge $Q^{\mathrm{u}p}$ remains constant, then decreases (at the merger) and starts to oscillate regularly once the initial relaxation period (which can be violent) is over. The process is similar to what we will show to happen in the quadrupolar case in Fig. 9 below.
@@
One can verify that the difference between the frequencies of the real and imaginary component of the scalar field is the origin of the dominant oscillation mode of the charge distribution. To see this, we make a spectral decomposition of the scalar field at a point outside the origin—at the origin, the imaginary part of $\phi$ is zero by construction. We see from Fig. 6 that the scalar field is practically monochromatic and the difference between the maximum of the real and imaginary parts gives $\omega_{R}-\omega_{I}=0.020$ (see the caption of Fig. 6 for more details), which is consistent with the dominant peak of $Q^{\rm up}$. In addition, we also find a second peak which is located around the first odd multiple of the base frequency, analogously to what was found in Xie et al. (2021) and consistent with the fact that a scalar potential has a $\mathbb{Z}_{2}$ symmetry. On the other hand, the small oscillations of $Q^{\rm up}$ that we already discussed corresponding the peak in the bottom right panel of Fig. 5, coincide very well with $\omega_{R}+\omega_{I}=1.898$ and can be shown to be produced by the next order perturbation term on the charge density $\propto j^{0}$, after the term with frequency $\omega_{R}-\omega_{I}$.
@@ III.2 Higher multipolar boson stars
We find that none of the first five types of spherical stars in Table 1 leads to charge-swapping configurations for the same range of $d$ as in the dipole case, not even extending up to a value of $d=48$. Black holes are often formed instead. In contrast, the last two types of stars with frequencies of $0.99$ and $0.995$ are capable of forming non-trivial configurations that are everywhere regular. Stationary boson stars with these frequencies are very dilute, as can be seen from their values of $\mathcal{C}$ and radius $R_{99}$ in Table 1. We have tried three $d$ distances for these two $\omega$’s respectively: 48, 42 and 36, and find that for both $\omega$’s no apparent horizon is formed for $d=42$ and 48, but for $d=36$ (and also smaller values), the initial clump collapses to a black hole. In Fig. 8, we can see the collapse of the lapse for one of the quadrupole cases together with the corresponding dipolar case which does not collapse.
@@
Configurations that avoid collapse exhibit quadrupolar charge-swapping structures. Despite differing shapes, these structures show quantitative similarities to the dipolar case, following similar correlations between initial parameters $\omega$ and $d$, and the final solution properties. However, their lifetimes differ markedly from self-interacting quadrupoles without gravity (i.e., quadrupolar Q-balls) Xie et al. (2021), with all cases persisting for at least $t=5000T_{0}$. In Fig. 9, snapshots of charge density illustrate typical behavior, comparing initial star distributions at $t=0$ with those within half a charge-swapping period later in the simulation. The bottom panel of Fig. 9 shows $Q^{\rm quad}$, the total charge associated with one blob by integrating $n_{\mu}j^{\mu}$ over the $z>|x|$ region.
In this case, compared to the analogous plot for a dipole shown in Fig. 4, the initial relaxation process begins later. Specifically, because the four stars are initially far apart, the Noether charge $Q^{\mathrm{quad}}$ remains positive until approximately $t/T_{0}\sim 50$, when the merger occurs. Following the merger, relaxation begins, and regular stable oscillations of the charge are observed after $t/T_{0}\sim 100$. We have found that the scalar field ejected at the merger of dipolar and quadrupolar cases is essential to obtain a charge-swapping configuration.
@@ IV.1 Sextic potential
until the turning point $\omega_{\rm min}$. This leads to one or two stable branches according to the value of $\alpha_{G}$; see Fig. 1. These observations are consistent with the conclusions drawn from a stability analysis of such solutions using catastrophe theory Tamaki and Sakai (2010); Kleihaus et al. (2012). After the turning point, all solutions are unstable regardless of the sign of $dM/d\omega$, some of them migrating toward a stable boson star and others dissipating to infinity or collapsing to a black hole depending on the binding energy, which is similar to the free field and the quartic self-interaction cases previously reported in the literature Seidel and Suen (1990); Balakrishna et al. (1998).
Compared to the free field case, the self-interacting solutions are qualitatively very different when the gravitational coupling is small, where they are close to their Q-ball counterparts. Nevertheless, these cases can give rise to very compact solutions (as $f(r=0)$ increases) that differ significantly from the Q-ball solutions and achieve very high compactness, making them different from both the mini-boson stars and the corresponding Q-balls; see Fig. 3. With this in mind, we have chosen the stars in Table 3 to prepare the charge-swapping configurations. In particular, we restrict our study to superpositions of single star solutions that are stable on their own. Of course, this does not guarantee that the complex boson star solution will be obtained, and neither ensures that the final solution will not be a black hole. Since we have fixed $\mathfrak{g}=1/2$ as our fiducial model, we only have two degrees of freedom to construct the initial data: $\omega$ and $d$.
@@
We shall explore the dipole superposition with initial separations $d=\{8,10,12,14,16\}$. Before proceeding with the boson star analysis, as a sanity check, we first re-produce the 3+1D results of charge-swapping Q-balls in Ref. Xie et al. (2021), The resulting flat spacetime evolutions are shown in Fig. 10. We will see that the charge-swapping configurations, when gravity is included, share many properties with the flat spacetime limit analogue. In this sense the Fig. 10 is representative of some of the properties that are obtained in general. There are four stages present when charge-swapping configurations are succesfully obtained: An initial relaxation stage, which ends around $t=500T_{0}$ for the case presented in Fig. 10; a plateau stage of charge-swapping, ending around $t=1000T_{0}$; a fast decay stage in which $Q^{\rm up}$ drops abruptly to zero and the energy density also decreases but towards a finite spherically symmetric state; a long-lived stage of the (highly perturbed) oscillaton Seidel and Suen (1991), the oscillon in the absence of gravity (see Ref. Zhang et al. (2022) for related real vector field solitons with cosmological applications).
Unlike the analysis conducted for the free field case, in all the scenarios explored with the sextic potential, we positioned the stars initially close together. For instance, in all the figures presented in this section, the maximum distance $d$ considered is 16. When comparing this to the radii of the stars listed in Table 3, it is evident that the stars are either “touching” or very close to each other initially in nearly all cases. Typically, the repulsive self-interaction from the quartic term is sufficient to prevent gravitational collapse, making the consideration of larger initial separations less relevant for the sextic potential compared to the free field scalar case. Consequently, the early stages of the relaxation process, including the merger, occur rapidly after the initial configuration of the stars.
We see from Fig. 10 that the longest lifetime is given by $\omega=0.8$ and $d=14$, consistent with that of Ref. Xie et al. (2021), among other results. These flat-space solutions have smaller radii than the free-field configurations with gravity discussed in Section III. Nevertheless, for an easy comparison, we keep the same numerical setting as in the dipolar free field case with gravity.
@@
Next, we switch on $\alpha_{G}$ to investigate the effect of introducing the gravitational interaction. To monitor the evolution in this case, although not perfect as it is now not really conserved, we will still use the $E$ integral as a proxy. In all the cases presented in this section, we observe that the envelope of the Noether charge in the upper hemisphere always follows a similar qualitative behavior to that of the $E(t)$ curve. In Fig. 11, we show the cases with frequency $\omega=0.8$ and gravitational coupling $0.0018\pi\alpha_{G}=$ 0.001 and $0.01$. This corresponds to the two first rows in Table 3. For $0.0018\pi\alpha_{G}=$ 0.001, $\omega=0.8$ is the point where the spherical boson stars start to deviate from the Q-ball counterparts (see Fig. 1). The same is also noticeable in the top panel of Fig. 11 for the charge-swapping configurations, where we see that the lifetimes are slightly shortened in the presence of gravity for all the $d$’s listed. For a smaller $\alpha_{G}$, the system basically reduces to the Q-ball case. For $0.018\pi\alpha_{G}=$ 0.01 in the bottom panel, interestingly, we see that the differences between the $E$ curves are negligible for different $d$’s, due to the significantly reduced lifetimes. This means that the gravitational interaction is already dominating for $0.018\pi\alpha_{G}=$ 0.01, and the stable end state corresponds to an oscillaton as expected. To confirm that the final stage is indeed an oscillaton, we examine the quantity $Q^{\mathrm{up}}$. In regions where this quantity approaches zero, we analyze the phase portraits of the complex scalar field, specifically plotting $\Re(\phi)$ versus $\Im(\phi)$. During the oscillaton stage, the phase portraits exhibit oscillations along a single linear direction.
@@
In these cases presented in Fig. 11, as well as in the $\alpha_{G}\to 0$ limit another relevant mechanism takes place in the formation of charge-swapping states. When self-interaction is significant this mechanism is characterized by a net attraction force arising from the nonlinear scalar field, which is specific to charge-swapping systems and has been detailed in Ref. Copeland et al. (2014) for Q-balls. In our configurations, when the quartic term of the scalar potential is large and the effective gravitational constant is therefore small, this property predominates in the attraction between the two blobs, facilitating the formation of bound states. In the free-field scenario, this effect is absent, with the gravitational interaction being the sole force that holds the stars together. As we will demonstrate, when we increase the value of $\alpha_{G}$ for the sextic potential case under examination in this section, the interaction that initially facilitated the formation of bound states becomes subdominant, allowing gravity to dominate the dynamics of the system.
@@
The evolution of $E$ for $\omega=0.75$ and $0.018\pi\alpha_{G}=$ 0.01 for three different $d$’s together with the total Noether charge in the $z>0$ region for $d=16$ is shown in Fig. 12. We can see how the charge is radiated away in the final stage. More interesting is the fact that for these cases, as well as for the cases in Fig. 13, the configuration transits through two charge-swapping plateaus with different $E$ and $Q^{\rm up}$ before migrating to the oscillaton configuration. This resembles the cascading of the energy for excited oscillons Wang et al. (2023). However, at the level of the quantities we have observed, a direct connection between the two phenomena has yet to be established.
Figure 12: Varying $d$ for fixed $\omega=0.75$. In the right panel we show the Noether charge in the $z>0$ region, $Q^{\rm up}$, only for the $d=12$ case. For reference we mark the different stages of the evolution for this case in both plots.
@@ IV.2 Logarithmic potential
As in the sextic potential case, here we also prepared configurations that are close to each other at $t=0$, so in this sense we are expected to start with states close to the charge-swapping configuration. Interestingly, similar to the charge-swapping Q-ball case with the logarithmic potential, there is no relaxation process that radiates away excess of the scalar field, as can be confirmed in Fig. 15. Also, the charge-swapping process begins very early in the simulations. The latter is reflected in the fact that the charge-swapping period $T_{\rm swap}$ is clearly defined from the beginning of the evolution, as can be seen in Fig. 16: $T_{\rm swap}$ is defined as twice the time elapsed between two contiguous zero points of the Noether charge in the upper half space $Q^{\rm up}$.
@@ V Anisotropic distribution of satellite galaxies
$(\Psi_{\rm w},U,x^{i},t)\rightarrow(\Lambda^{2}\Psi_{\rm w},\Lambda^{2}U,\Lambda^{-1}x^{i},\Lambda^{-2}t)\,,$ (69)
for any value of the real scaling constant $\Lambda$. This allows to obtain the full sequence of Newtonian boson stars once a single solution has been found, unlike the relativistic analogue, where the family of solutions, say, the one presented in Fig. 1, must be constructed numerically solving the equations at each point. This rescaling is independent of the rescaling used to obtain dimensionless quantities, so to recover the physical system, as in the relativistic case, we must choose a value of $\mu$ as explained in Section II.3. We have previously obtained the metrics in the full general relativistic formulation. For this application, it is sufficient to extract the non-relativistic limit from Eq. (67) and Eq. (68).
We choose to perform the analysis at a certain representative space slice with $t=4000T_{0}$ of the dipolar configuration with $\omega=0.98$ well within the charge-swapping stage. However, the results are insensitive to the specific time period chosen within this stage. To extract the 2D data at such time we use the Kuibit tool Bozzola (2021). Based on the consistency statistical tests performed in Solís-López et al. (2021), we choose the values of $\mu$ and $\Lambda$, which completely fix the scale of the system, to be such that the size of the configuration, which has an equatorial radius of about 16 units, corresponds to a physical size of 300 kpc (radius encompassing the 11 classical satellites of the Milky Way Pawlowski and Kroupa (2020)), together with the condition that the circular velocity of the stars in the galactic disk at a point 30kpc away from the galactic center is 100 km/s, implying similar contributions from the enclosed mass at radius r=30kpc of dark matter and baryonic matter. These two conditions lead to the values $\mu\hbar\sim 10^{-25}\text{eV}$ and $\Lambda=0.01$. The precise determination of these two parameters given the rotation curve of the Milky Way and a model for the galactic bulge and disc deserves a statistical analysis which is beyond the scope of this paper.
@@
Figure 19: Angular distribution of the angular momenta of $10^{4}$ “particles” for the fiducial dipolar charge-swapping configuration. The initial distribution is isotropic, and it quickly relaxes to a configuration where the tangent of the orbital poles are oriented toward the galactic plane $\theta=\pi/2$ We only include particles within $\Lambda\mu R\in(0.16,16)$ which takes into account “particles” with $R>30$ kpc, the size of the Milky Way stellar disc.
@@
after restoring units. Meaning that the angular distribution of the angular momenta is already oriented towar the galactic pole after 1 Gyr (as can be seen in Fig. 19) and that the charge-swapping configurations live for much longer than is required for the galactic halos in the Universe.
@@
Figure 21: “Particles” in a disc of radius $30$ kpc and thickness $3$ kpc moving in a (Newtonian) gravitational potential of the fiducial dipolar configuration (see Fig. 18). The side of the cube is of length $60$ kpc. The bottom panels display a top-view of the disc.